solve the following: substitution method 3x-2y=-7 y=2x+6
Answer by checkley77(12844)

3x-2y=-7
y=2x+6
3x-2(2x+6)=-7
3x-4x-12=-7
3x-4x=-7+12
-x=5
x=-5 ans.
3*-5-2y=-7
-15-2y=-7
-2y=-7+15
-2y=8
y=8/-2
y=-4 ans.